A person's blood glucose level and diabetes are closely related. Let x be a random variable measured in milligrams of glucose per deciliter (1/10 of a liter) of blood. Suppose that after a 12-hour fast, the random variable x will have a distribution that is approximately normal with mean μ = 83 and standard deviation σ = 22. Note: After 50 years of age, both the mean and standard deviation tend to increase. For an adult (under 50) after a 12-hour fast, find the following probabilities. (Round your answers to four decimal places.)
(a) x is more than 60

(b) x is less than 110

(c) x is between 60 and 110

(d) x is greater than 140 (borderline diabetes starts at 140)

Answer by stanbon(75887)

A person's blood glucose level and diabetes are closely related.
Let x be a random variable measured in milligrams of glucose per deciliter (1/10 of a liter) of blood.
Suppose that after a 12-hour fast, the random variable x will have a distribution that is approximately normal with mean μ = 83 and standard deviation σ = 22. Note: After 50 years of age, both the mean and standard deviation tend to increase. For an adult (under 50) after a 12-hour fast, find the following probabilities. (Round your answers to four decimal places.)
(a) x is more than 60
z(60) = (60-83)/22 = -1.0455
P(x > 60) = P(z > -1.0455) = 0.8521
------------------------------------------
(b) x is less than 110
z(110) = (110-83)/22 = 1.2273
P(x < 110) = P(z < 1.2273) = 0.8901
-------------------------------------------
(c) x is between 60 and 110
Find the 2 z-values; then find the area between them.
--------------------
(d) x is greater than 140 (borderline diabetes starts at 140)
Find the z-value; then find its right tail.
==================================================
Cheers,
Stan H.